The longest escalator in the world is located in Hong Kong, China. The escalator has a length of 745 ft and rises 377 ft vertical from start to finish. What is the horizontal distance from the bottom of the escalator to the top of the escalator?
A) 368 B) 643 C) 412896 D) 377 E) 600

Answer by checkley77(12844)

745^2+377^2=x^2
555,025+142,129=x^2
x^2=697,154
x=sqrt697,154
x=834.957 ans.